how much of a 50% saline solution must be added to 100ml of 20% saline solution to make a mixture that is 40% saline?

Answer by stanbon(75887)

how much of a 50% saline solution must be added to 100ml of 20% saline solution to make a mixture that is 40% saline?
---------------------------------
Equation:
active + active = active
0.20(100) + 0.50x = 0.40(100+x)
Multiply thru by 100 and solve for "x":
20*100 + 50x = 40*100 + 40x
10x = 20*100
x = 200 ml (amt of 50% solution needed for the mixtue)
=========================================================
Cheers,
Stan H.
===============================================